OpenStudy (anonymous):

Quick check? Thank You for helping! Use the formula v1*v2=|v1||v2|cos theta to find the cosine of the angle between the two vectors (-7,8) and (5,3). I got cos theta=-11/61.97. Is this correct? Thanks!

OpenStudy (anonymous):

@nincompoop @Michele_Laino @ganeshie8 @campbell_st @zepdrix @Luigi0210 @Destinymasha @radar @Whitemonsterbunny17

OpenStudy (anonymous):

I also solved for theta, getting 100.22 degrees.

OpenStudy (michele_laino):

we can write this: \[\left\| {{v_1}} \right\| = \sqrt {49 + 64} = 10.63\] \[\left\| {{v_2}} \right\| = \sqrt {25 + 9} = 5.83\] \[{v_1} \cdot {v_2} = - 35 + 24 = - 11\] so we get: \[\cos \theta = \frac{{ - 11}}{{10.63 \times 5.83}}\]

OpenStudy (anonymous):

so same thing, awesome!

OpenStudy (anonymous):

Thanks for the confirmation!

OpenStudy (michele_laino):

yes! \[\cos \theta = \frac{{ - 11}}{{10.63 \times 5.83}} = - \frac{{11}}{{61.97}}\]

OpenStudy (michele_laino):

:)